# Seeing Straight: Document Orientation Detection for Efficient OCR

### v1

@@ Abstract
Despite significant advances in document understanding, determining the correct orientation of scanned or photographed documents remains a critical pre-processing step in the real world settings. Accurate rotation correction is essential for enhancing the performance of downstream tasks such as Optical Character Recognition (OCR) where misalignment commonly arises due to user errors, particularly incorrect base orientations of the camera during capture. In this study, we first introduce OCR-Rotation-Bench (ORB), a new benchmark for evaluating OCR robustness to image rotations, comprising (i) ORB-En, built from rotation-transformed structured and free-form English OCR datasets, and (ii) ORB-Indic, a novel multilingual set spanning 11 Indic mid to low-resource languages. We also present a fast, robust and lightweight rotation classification pipeline built on the vision encoder of Phi-3.5-Vision model with dynamic image cropping, fine-tuned specifically for 4-class rotation task in a standalone fashion. Our method achieves near-perfect 96% and 92% accuracy on identifying the rotations respectively on both the datasets. Beyond classification, we demonstrate the critical role of our module in boosting OCR performance: closed-source (up to 14%) and open-weights models (up to 4x) in the simulated real-world setting.
Seeing Straight: Document Orientation Detection for Efficient OCR
Suranjan Goswami1, Abhinav Ravi2, Raja Kolla2, Ali Faraz2, Shaharukh Khan2 Akash1, Chandra Khatri2 and Shubham Agarwal2 1OLA Electric, Bangalore, India 2Krutrim AI, Bangalore, India {suranjan.goswami, akash.shyam}@olaelectric.com {raja.kolla, ali.faraz, shubham.agarwal1}@olakrutrim.com
@@ 1 Introduction
Recent advancements in Vision Language Models (VLMs) bai2023qwenvl; chen2024far; lu2024deepseek-vl; beyer2024paligemma; abdin2024phi2; xue2024xgen-mm; xiao2024florence have demonstrated impressive capabilities in general purpose visual-linguistic understanding. Models such as DocOwl series (hu2024mplug; docowl2_2024), DocLLM (wang2023docllm) and Donut kim2022donut have significantly advanced document understanding tasks.
@@
Even the benchmark evaluations are typically conducted on curated datasets like RVL-CDIP harley2015icdar, FUNSD jaume2019funsd, and DocVQA mathew2021docvqa, inter alia, where documents are upright and well aligned. Orientation is usually assumed or corrected as a pre-processing step, limiting their robustness in practical scenarios. In contrast, real-world documents such as receipts, forms, handwritten notes, and IDs are often captured via mobile phones or scanners under uncontrolled conditions, resulting in unintended rotations. These misalignments can significantly degrade OCR and information extraction accuracy. Furthermore, existing evaluation datasets are heavily skewed toward English or Latin-script documents, with limited insights for multilingual settings, especially for Indian languages. Resources like Wikisource [^1] https://wikisource.org/, an open-access repository of digitized literary works, present a promising avenue for constructing large-scale multilingual OCR benchmarks. However, the lack of standardized, rotation-diverse, and multilingual datasets remains a bottleneck in evaluating and improving generalization of current OCR engines.
To address these gaps, we propose a lightweight rotation classification module that operates as a pre-processing step to OCR in noisy, real-world settings. Specifically, we fine-tune the vision encoder of Phi-3.5-Vision-Instruct abdin2024phi3 as a standalone module for this task, improving OCR robustness. We formulate orientation correction as a four-class classification task over canonical angles, following prior works  (unnikrishnan2009combined), effective for most documents which are typically scanned or captured in vertical or horizontal orientations. Our method is efficient, accurate, and generalizes across languages and scripts, enabling reliable alignment of rotated document images before text recognition.
In addition, we further introduce a novel multilingual OCR benchmark spanning 11 mid- to low-resource Indic languages, constructed from high-quality document-text pairs sourced from Wikisource. To complement this, we incorporate two standard English OCR datasets, transformed with controlled rotations to simulate real-world misalignment scenarios. Together, these form a comprehensive benchmark for evaluating OCR robustness to rotation. We evaluate our rotation correction module alongside state-of-the-art OCR models and demonstrate consistent improvements in performance across both English and Indic scripts. Our key contributions could be summarized as:
@@
  • We present OCR-Rotation-Bench (ORB), a novel benchmark for evaluating OCR robustness to practical image rotation scenarios. ORB consists of two subsets: (i) ORB-En, derived from rotated versions of 2 standard English datasets, and (ii) ORB-Indic, a curated set of 328 rotated real-world document images with OCR ground truth (GT) annotation across 11 Indic languages.
  • We conduct an exhaustive evaluation of downstream OCR performance, demonstrating significant gains in accuracy when a correction module is integrated.
  • We will publicly release our trained models and evaluation datasets and code to support reproducibility and community benchmarking.
@@ 2 Related Work
Rotation Correction in Document Pre-processing: Orientation correction is a fundamental step in OCR pipelines, particularly for documents captured via mobile devices or scanners. Classical methods such as projection profiles, Hough transforms, and connected component analysis estimate dominant text skew using low-level visual features  bao2022adaptive. al. bao2022adaptive. While effective on clean, structured inputs, these approaches degrade under noise, script diversity, or low resolution.
unnikrishnan2009combined formalized rotation into four canonical classes: $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ ($90^{\circ}$ in our work for clarity)  hu2001automatic explored symmetry-based textual analysis in natural scenes, illustrating the effectiveness of spatial text patterns for orientation detection. While recent deep learning based approaches have explored implicit orientation correction  zhao2022transformer, they remain limited to English language. In contrast, our work isolates and enhances the rotation step, enabling language-agnostic benchmarking and seamless integration into existing OCR workflows.
@@
Optical Character Recognition (OCR): OCR systems have progressed from traditional rule-based engines like Tesseract smith2007overview to deep learning models that integrate visual and sequential components. Modern approaches such as docTR liao2023doctr, and TrOCR li2021trocr demonstrate strong performance on both printed and handwritten text in English, leveraging transformer-based visual encoders and pretrained language decoders. However, these models typically assume upright inputs, and even minor rotations can significantly degrade recognition accuracy, highlighting the need for explicit orientation correction. On the other hand, recent multilingual OCR systems, such as the closed-source Chitrapathak, GPT-4o, Gemini-2.5 Flash and Gemini-2.5 Pro222 https://ai-labs.olakrutrim.com/ https://platform.openai.com/docs/models/gpt-4o/ https://deepmind.google/models/gemini/ have broadened language support. Recently released open-source solution like Nanonets nanonets2024ocr claims to provide industrial-grade OCR pipelines, utilizing fine-tuned Qwen-VL backbones qwen2023vl, effectively bridging foundational vision-language models with document-specific applications in the multi-lingual setting. We benchmark the performance of these models on both structured and free form generations with our trained module.
@@
VLMs for Document Understanding: Recent VLMs such as the LLaVA family (liu2024visual; liu2024improved) have advanced multimodal alignment by coupling vision encoders (radford2021learning; zhai2023sigmoid; tschannen2025siglip) with large language models via visual instruction tuning. Several successors (lu2024deepseek; laurenccon2024matters; tong2024cambrian; xue2024xgen; team2024gemma) follow similar paradigms, excelling in translation, captioning, and visual reasoning. Document-centric variants such as DocOwl (hu2024mplug; docowl2_2024), DocLLM (wang2023docllm), and Donut (kim2022donut) with emerging datasets like BigDocs (bigdocs2024), show promise but remain English-centric and rotation-agnostic.
@@
Support for Indic languages in VLMs remains limited. Models such as Chitrarth (khan2025chitrarth) and Chitranuvad (khan2025chitranuvad), built on Krutrim (kallappa2025krutrim), represent early efforts to support 10 Indian languages and low-resource translation respectively. Others, including PALO (maaz2024palo), Maya (alam2025behind), and Pangea (yue2024pangea), partially supports select languages like Hindi, Bengali, Tamil, and Telugu but remain general-purpose VLMs. However, none explicitly address document understanding or orientation robustness in low-resource Indian contexts which our work directly addresses.
@@ 3 Model Architecture
We define the architecture of our rotation classifier model in the following. Let $x\in\mathbb{R}^{H\times W\times 3}$ denote an input RGB image and the corresponding ground-truth rotation label belongs to the set of classes $\mathcal{Y}=\{0,1,2,3\}$, where 0 corresponds to $90^{\circ}$, 1 to $0^{\circ}$, 2 to $90^{\circ}$, and 3 to $180^{\circ}$. To enhance robustness and model generalization, we adopt a dynamic cropping strategy  dong2024internvl, where each input image $x$ is transformed into multiple spatial crops $\{x^{(1)},x^{(2)},\dots,x^{(K)}\}$. Each crop is independently passed through the vision encoder to produce a sequence of visual tokens $X^{(k)}\in\mathbb{R}^{L\times D}$ where $D$ denotes embedding dimension, $L$ and $K$ the number of visual tokens and crops respectively. The vision encoder consists of a patch embedding module followed by a transformer encoder, consistent with the modern VLM architectures.
From each crop’s output sequence, we extract the CLS token from the first position, $h_{\text{cls}}^{(k)}=X_{0}^{(k)}\in\mathbb{R}^{D}$, which captures global contextual information, allowing to aggregate spatial and semantic cues for that crop.
We compute the average CLS embedding over all $K$ crops of the image to obtain a robust, unified image representation.
We then use a classification head, which is a lightweight feedforward neural network $f_{\text{clf}}$ that maps $h_{\text{cls}}$ to one of four rotation classes:
@@
where $W_{1}\in\mathbb{R}^{D^{\prime}\times D}$ and $b_{1}\in\mathbb{R}^{D^{\prime}}$ denote the weights and bias of the first layer, while $W_{2}\in\mathbb{R}^{C\times D^{\prime}}$ and $b_{2}\in\mathbb{R}^{C}$ represent those of the output layer. We use the GELU activation (hendrycks2016gelu) function $\phi$, with $C=4$ denoting the number of rotation classes, and $D^{\prime}=\frac{D}{2}$ specifying the hidden size of the intermediate layer. To mitigate overfitting, we applied dropout (srivastava2014dropout) as $h_{\text{drop}}=\text{Dropout}(h_{\text{cls}})$ with a probability of 0.2.
Thus, the complete forward pass of the classification head is:
@@
$\hat{y}$ $=W_{2}\cdot z+b_{2}\in\mathbb{R}^{4}$
The predicted logits $\hat{y}$ are trained using softmax cross-entropy loss. This architecture enables compact and efficient inference while maintaining high classification accuracy. We adapt the vision encoder from the Phi-3.5 model abdin2024phi3 as the backbone for our rotation classification task. By fine-tuning the pretrained vision encoder standalone on our proprietary document image datasets, we obtain a lightweight yet semantically powerful model that leverages the representational strength of CLIP ViT-L/14 model while being optimized for rotation detection in our pipeline.
@@ 4 OCR Rotation Benchmark (ORB)
To address the lack of standardized benchmarks for rotation correction in OCR pipelines, we introduce OCR-Rotation-Bench (ORB), a new benchmark for evaluating OCR robustness to image rotations. ORB comprises two subsets: one for English-language documents (ORB-En) and another for multilingual Indic scripts (ORB-Indic). Both the datasets are evenly distributed across four canonical rotation angles,
enabling balanced evaluation without any orientation bias.
ORB-En:
We select two diverse and publicly available datasets to enable
  • SROIE 2019 huang2019sroie, a widely-used benchmark comprising real-world scanned documents such as receipts and invoices, annotated at the field level for structured OCR tasks.
  • SynthDog kim2022donut, on the other hand, contains free-form synthetically generated English documents with complete OCR ground truth, offering controlled document structure and high annotation fidelity.
We rotate the test and validation sets containing 347 and 500 images, respectively of the two datasets uniformly across the four rotation classes to simulate real-world scenarios such as upside-down or sideways documents often seen in mobile or bulk-scanned OCR settings. This combination of real and synthetic data ensures that the benchmark is both practically grounded and generalizable across OCR use cases.
Multilingual Image Rotation Classification GT OCR Output
Refer to caption $90^{\circ}$ Refer to caption
Refer to caption $0^{\circ}$ Refer to caption
Refer to caption $90^{\circ}$ Refer to caption
Refer to caption $180^{\circ}$ Refer to caption
Figure 2: Sample images from the ORB-Indic dataset along with their predicted rotation classes and corresponding ground truth OCR outputs. Each row presents an input document image, the predicted orientation label, and the correct OCR result. The rotation classification model accurately predicted the orientation for all examples shown.
ORB-Indic: To construct the multilingual benchmark for rotation-aware OCR, we curate a high-quality Indic-language dataset from Wikisource, an open-access repository of digitized literary works. The dataset spans 11 Indic languages: Hindi, Bengali, Tamil, Malayalam, Telugu, Marathi, Kannada, Gujarati, Punjabi, Oriya, and Assamese, covering a wide typographic and linguistic diversity representative of real-world document understanding. Table 1 provides the balanced distribution of samples in each language. However, unlike concurrent work (anonymous2025multimodal), our dataset additionally includes Assamese and specifically emphasizes robustness to rotation variations. We develop custom web crawlers to extract both document images and corresponding OCR text based on structured metadata provided by Wikisource. Specifically,
@@
  • The corresponding OCR text is extracted via the pagetext tag, which reflects the latest verified human-edited transcription.
To further ensure the quality and alignment of image-text pairs, we manually clean the collected data, discarding samples with mismatches, formatting issues, or incomplete annotations. The final test set comprises 328 document images, evenly distributed across the 11 languages and four canonical rotation classes. To the best of our knowledge, ORB-Indic is the first multilingual benchmark explicitly designed to evaluate the effect of rotation on OCR performance in Indic scripts. It provides a valuable resource for benchmarking and advancing rotation-robust OCR models in low-resource and multilingual settings.

### v2

@@ Abstract
Despite significant advances in document understanding, determining the correct orientation of scanned or photographed documents remains a critical pre-processing step in the real world settings. Accurate rotation correction is essential for enhancing the performance of downstream tasks such as Optical Character Recognition (OCR) where misalignment commonly arises due to user errors, particularly incorrect base orientations of the camera during capture. In this study, we first introduce OCR-Rotation-Bench (ORB), a new benchmark consisting of 1863 images for evaluating OCR robustness to rotations, comprising (i) ORB-En, built from rotation-transformed structured and free-form English OCR datasets, and (ii) ORB-Indic, a novel multilingual set spanning 11 Indic mid to low-resource languages. We also present a fast, robust and lightweight rotation classification pipeline built on the vision encoder of Phi-3.5-Vision model with dynamic image cropping, fine-tuned specifically for 12-class rotation task in a standalone fashion. Our method achieves near-perfect 98% and 96% accuracy on identifying the rotations respectively on both the datasets. Beyond classification, we demonstrate the critical role of our module in boosting OCR performance: closed-source (up to 20%) and open-weights models (up to 4x) in the simulated real-world settings.
@@ 1 Introduction
Recent advancements in Vision Language Models (VLMs) [5, 10, 38, 8, 2, 61, 59] have demonstrated impressive capabilities in general purpose visual-linguistic understanding. Models such as DocOwl series [20, 21], DocLLM [57] and Donut [29] have significantly advanced document understanding tasks.
@@
Even the benchmark evaluations are typically conducted on curated datasets like RVL-CDIP [18], FUNSD [25], and DocVQA [40], inter alia, where documents are upright and well aligned. Orientation is usually assumed or corrected as a pre-processing step, limiting their robustness in practical scenarios. In contrast, real-world documents such as receipts, forms, handwritten notes, and IDs are often captured via mobile phones or scanners under uncontrolled conditions, resulting in unintended rotations. These misalignments can significantly degrade OCR and information extraction accuracy. Furthermore, existing evaluation datasets are heavily skewed toward English or Latin-script documents, with limited insights for multilingual settings, especially for Indian languages. Resources like Wikisource [^1] https://wikisource.org/, an open-access repository of digitized literary works, present a promising avenue for constructing large-scale multilingual OCR benchmarks. However, the lack of standardized, rotation-diverse, and multilingual datasets remains a bottleneck in evaluating and improving generalization of current OCR engines.
To address these gaps, we propose a lightweight rotation classification module that operates as a pre-processing step to OCR in noisy, real-world settings. Specifically, we fine-tune the vision encoder of Phi-3.5-Vision-Instruct [1] as a standalone module for this task, improving OCR robustness. We formulate orientation correction as a twelve-class classification task over 30 degree angle variations, extending prior works  [56], effective for most documents which are typically scanned or captured in vertical or horizontal orientations. Our method is efficient, accurate, and generalizes across languages and scripts, enabling reliable alignment of rotated document images before text recognition.
In addition, we further introduce a novel multilingual OCR benchmark spanning 11 mid- to low-resource Indic languages, constructed from high-quality document-text pairs sourced from Wikisource. To complement this, we incorporate three standard English OCR datasets, transformed with controlled rotations to simulate real-world misalignment scenarios. Together, these form a comprehensive benchmark for evaluating OCR robustness to rotation. We evaluate our rotation correction module alongside state-of-the-art OCR models and demonstrate consistent improvements in performance across both English and Indic scripts. Our key contributions could be summarized as:
  • We present OCR-Rotation-Bench (ORB), a novel benchmark for evaluating OCR robustness to practical image rotation scenarios. ORB consists of two subsets: (i) ORB-En, derived from rotated versions of 3 standard English OCR datasets of total 897 images, and (ii) ORB-Indic, a curated set of 966 rotated real-world multilingual document images with OCR ground truth (GT) annotation across 11 Indic languages.
@@
  • We also conduct an exhaustive evaluation of downstream OCR performance, demonstrating significant gains in accuracy when a correction module is integrated. We will publicly release all the related artifacts to support reproducibility and community benchmarking.
@@ 2 Related Work
Rotation Correction in Document Pre-processing: Orientation correction is a fundamental step in OCR pipelines, particularly for documents captured via mobile devices or scanners. Classical methods such as projection profiles, Hough transforms, and connected component analysis estimate dominant text skew using low-level visual features  [7]. While effective on clean, structured inputs, these approaches degrade under noise, script diversity, or low resolution.  Unnikrishnan and Smith [56] formalized rotation into four canonical classes: $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$.  Hu et al. [22] explored symmetry-based textual analysis in natural scenes, illustrating the effectiveness of spatial patterns for orientation detection. While recent deep learning approaches explore implicit orientation correction  [65], they remain limited to English. In contrast, our work isolates and enhances the rotation step, enabling language-agnostic benchmarking and seamless integration into existing OCR workflows.
@@
Optical Character Recognition (OCR): OCR systems have progressed from traditional rule-based engines like Tesseract [49] to deep learning models that integrate visual and sequential components. Modern approaches such as docTR [34], and TrOCR [33] demonstrate strong performance on both printed and handwritten text in English, leveraging transformer-based visual encoders and pretrained language decoders. However, these models typically assume upright inputs, and even minor rotations can significantly degrade recognition accuracy, highlighting the need for explicit orientation correction. On the other hand, recent multilingual OCR systems, such as the closed-source Chitrapathak, GPT-4o, Gemini-2.5 Flash and Gemini-2.5 Pro222 https://ai-labs.olakrutrim.com/ https://platform.openai.com/docs/models/gpt-4o/ https://deepmind.google/models/gemini/ have broadened language support. Recently released open-source solution like Nanonets [41] claims to provide industrial-grade OCR pipelines, utilizing fine-tuned Qwen-VL backbones [6], effectively bridging foundational vision-language models with document-specific applications in the multi-lingual setting. We show the performance of these models on both structured and free form generations with our trained module.
@@
VLMs for Document Understanding: Recent VLMs such as the LLaVA family [36, 35] have advanced multimodal alignment by coupling vision encoders [45, 63, 55] with large language models via visual instruction tuning. Several successors [37, 31, 53, 60, 52] follow similar paradigms, excelling in translation, captioning, and visual reasoning. Document-centric variants such as DocOwl [20, 21], DocLLM [57], and Donut [29] with emerging datasets like BigDocs [48], show promise but remain English-centric and rotation-agnostic.
@@
Support for Indic languages in VLMs remains limited. Models such as Chitrarth [28] and Chitranuvad [27], built on Krutrim [26], represent early efforts to support 10 Indian languages and low-resource translation respectively. Others, including PALO [39], Maya [3], and Pangea [62], partially supports select languages like Hindi, Bengali, Tamil, and Telugu but remain general-purpose VLMs. However, none explicitly address document understanding or orientation robustness in low-resource Indian contexts, which our work directly addresses.
@@ 3 Model Architecture
We define the architecture of our rotation classifier model as the following. Let $x\in\mathbb{R}^{H\times W\times 3}$ denote an input RGB image and the corresponding ground-truth rotation label belongs to the set of classes $\mathcal{Y}=\{0,1,2,3...,11\}$, where 0 corresponds to $0^{\circ}$, 1 to $30^{\circ}$, 2 to $60^{\circ}$, and so on, till $330^{\circ}$. To enhance robustness and model generalization, we adopt a dynamic cropping strategy  [14], where each input image $x$ is transformed into multiple spatial crops $\{x^{(1)},x^{(2)},\dots,x^{(K)}\}$. Each crop is independently passed through the vision encoder to produce a sequence of visual tokens $X^{(k)}\in\mathbb{R}^{L\times D}$ where $D$ denotes embedding dimension, $L$ and $K$ the number of visual tokens and crops respectively. The vision encoder consists of a patch embedding module followed by a transformer encoder, consistent with the modern VLM architectures. From each crop’s output sequence, we extract the CLS token from the first position, $h_{\text{cls}}^{(k)}=X_{0}^{(k)}\in\mathbb{R}^{D}$, which captures global contextual information, allowing to aggregate spatial and semantic cues for that crop. We compute the average CLS embedding over all $K$ crops of the image to obtain a robust, unified image representation. We then use a classification head, which is a lightweight feedforward neural network $f_{\text{clf}}$ that maps $h_{\text{cls}}$ to one of 12 rotation classes:
@@
where $W_{1}\in\mathbb{R}^{D^{\prime}\times D}$ and $b_{1}\in\mathbb{R}^{D^{\prime}}$ denote the weights and bias of the first layer, while $W_{2}\in\mathbb{R}^{C\times D^{\prime}}$ and $b_{2}\in\mathbb{R}^{C}$ represent those of the output layer. We use the GELU activation [19] function $\phi$, with $C$ denoting the number of rotation classes, and $D^{\prime}=\frac{D}{2}$ specifying the hidden size of the intermediate layer. To mitigate overfitting, we applied dropout [50] as $h_{\text{drop}}=\text{Dropout}(h_{\text{cls}})$ with a probability of 0.2. Thus, the complete forward pass of the classification head is:
@@
$\hat{y}$ $=W_{2}\cdot z+b_{2}\in\mathbb{R}^{C}$
The predicted logits $\hat{y}$ are trained using softmax cross-entropy loss. For the 4-class setting, we set $C=4$, while for fine-grained rotation classificaiton we use $C=12$, without modifying the rest of the pipeline. This architecture enables compact and efficient inference while maintaining high classification accuracy. We adapt the vision encoder from the Phi-3.5 model [1] as the backbone for our rotation classification task. By fine-tuning the pretrained vision encoder standalone on our proprietary document image datasets, we obtain a lightweight yet semantically powerful model that leverages the representational strength of CLIP ViT-L/14 model while being optimized for rotation detection.
@@ 4 OCR Rotation Benchmark (ORB)
To address the lack of standardized benchmarks for rotation correction in OCR pipelines, we introduce OCR-Rotation-Bench (ORB), a new benchmark for evaluating OCR robustness to image rotations. ORB comprises two subsets: one for English-language documents (ORB-En) and another for multilingual Indic scripts (ORB-Indic). Both the datasets are evenly distributed across either four or twelve rotation angles, enabling balanced evaluation without any orientation bias.
ORB-En: We select three diverse and publicly available datasets to enable standardized, reproducible evaluation in the English OCR tasks:
  • SROIE 2019 [23], a widely-used benchmark comprising real-world scanned documents such as receipts and invoices, annotated at the field level for structured OCR tasks.
  • SynthDog [29], which contains free-form synthetically generated English documents with complete OCR ground truth, offering controlled document structure and high annotation fidelity.
  • FUNSD [25], a form understanding benchmark consisting of scanned forms with word-level text annotations, semantic entity labels, and explicit link annotations, enabling joint evaluation of text recognition and document layout understanding.
We rotate the test and validation sets containing 347, 500 and 50 images, respectively of the three datasets uniformly across the 4 or 12 rotation classes to simulate real-world scenarios such as upside-down or angled documents often seen in bulk-scanned OCR settings. This combination of real and synthetic data ensures that the benchmark is practically grounded and generalizable across OCR use-cases.
Multilingual Image Rotation Classification English Image Rotation Classification
Refer to caption $270^{\circ}$ Refer to caption $0^{\circ}$
Refer to caption $150^{\circ}$ Refer to caption $60^{\circ}$
Figure 2: Qualitative evaluation of rotation classification. Multilingual ORB-Indic samples and real-world documents are shown together, illustrating accurate orientation prediction and successful downstream OCR for both English and Indic languages.
ORB-Indic: To construct the multilingual benchmark for rotation-aware OCR, we curate a high-quality Indic-language dataset from Wikisource, an open-access repository of digitized literary works. The dataset spans 11 Indic languages: Hindi, Bengali, Tamil, Malayalam, Telugu, Marathi, Kannada, Gujarati, Punjabi, Oriya, and Assamese, covering a wide typographic and linguistic diversity representative of real-world document understanding. Table 1 provides the balanced distribution of samples in each language. However, unlike concurrent work [4], our dataset additionally includes Assamese and specifically emphasizes robustness to rotation variations. We develop custom web crawlers to extract both document images and corresponding OCR text based on structured metadata provided by Wikisource. Specifically,
@@
  • The corresponding OCR text is extracted via the pagetext tag, which reflects the best human-edited transcription.
To further ensure the quality and alignment of image-text pairs, we manually clean the collected data, discarding samples with mismatches, formatting issues, or incomplete annotations. The final test set comprises 966 document images, distributed across the 11 languages and twelve rotation classes. To the best of our knowledge, ORB-Indic is the first multilingual benchmark explicitly designed to evaluate the effect of rotation on OCR performance in Indic scripts. It provides a valuable resource for benchmarking and advancing rotation-robust OCR models in low-resource and multilingual settings.